# Zagreb, Croatia Ensembling uncertainty with negative objectness

### v3

@@ 4 Training UNO with and without real negative data
Every training minibatch contains an equal number of samples for all K+1 classes. Thus, we expose our model to significantly less negative data than alternative approaches [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li, Yu and Aizawa(2019), Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu, Hendrycks and Gimpel(2017)] that utilize 2×2$\times$2 × more negatives than inliers. Even though the outlier exposure of our model is weak, it will still be biased towards detection of the test outliers that are similar to the training negatives. To circumvent this issue, we replace the auxiliary negative dataset with synthetic negatives produced by a generative model [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcić et al.(2021)Grcić, Bevandić, and Šegvić].
Synthetic negatives. We generate model-specific synthetic negatives by a normalizing flow $f_{\psi}$ [Dinh et al.(2017)Dinh, Sohl-Dickstein, and Bengio, Kingma and Dhariwal(2018)] that we train alongside our discriminative classifier $f_{\theta}$. We train the flow to generate samples that resemble inliers, but at the same time give raise to uncertain predictions across the inlier classes according to the following loss [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic]:
@@
The first loss term $L_{\mathrm{mle}}(\psi)=\textbf{E}_{\mathbf{x^{+}}\in D_{\text{in}}}[-\ln p_{\psi}(\mathbf{x})]$ maximizes the likelihood of inlier samples, while the second loss term $L_{\text{jsd}}(\psi,\theta)=\text{JSD}(U,f_{\theta}(\mathbf{x}))$ minimizes the JS divergence between uniform distribution and the predictions over the K inlier classes. The two competing objectives settle down when the generative model produces samples at the border of the inlier manifold [Lee et al.(2018)Lee, Lee, Lee, and Shin]. The above loss has to jointly optimize $f_{\theta}$ and $f_{\psi}$ so that the discriminative model gets aware of the negative class and thus preclude feature collapse [Lucas et al.(2019)Lucas, Shmelkov, Alahari, Schmid, and Verbeek]. Of course, the training also has to ensure that $f_{\theta}$ satisfies the primary discriminative loss.
However, a naive application of this recipe in the K+1-way context would not work as intended. Infact, the generative loss (6) and the discriminative loss (5) would simply be satisfied by confident classification of all generated samples into the negative class. This trivial solution would embed the generated samples into a compact region of the feature space, and make them useless as proxies for outlier detection. Thus, we propose a two-step optimization procedure. The first step jointly trains $f_{\psi}$, $f_{\theta}$, and the K inlier logits according to (6). The second step freezes the flow and fine-tunes the loss (5) in order to ensure recognition of generated negatives as the K+1-th class. Different from previous approaches [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic], the loss (6) serves only to promote generation of useful synthetic negatives, while we ensure the outlier recognition quality by subsequent fine-tuning according to the loss (5).
@@
Figure 4 shows the scaling of our training approach to dense recognition. We paste negative training content atop the regular image to produce a mixed-content image. The mixed-content image is fed to K+1-way classifier that optimizes the cross-entropy objective, similar to the image-wide setup. In the case of synthetic negatives, we jointly optimize the dense classifier and the normalizing flow.
Figure 4: Fine-tuning of dense classifier equipped with UNO. We paste negative training data (either real or synthetic) atop regular inlier images. The resulting mixed-content image is fed to the dense classifier that optimizes cross-entropy loss over K+1 classes.
@@ 5 Experiments
Experimental setup. We evaluate the performance of UNO on standard image-wide and pixel-level outlier detection benchmarks. In the pixel-level setup, we consider Fishyscapes [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] and SMIYC [Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann], the two prominent benchmarks for semantic anomaly segmentation in road driving scenes. In the case of image-wide setup, we consider the CIFAR-10 and ImageNet-200 setups from the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. We report the standard evaluation metrics AUROC and FPR95 in image-wide, and AP and FPR95 in dense prediction experiments. Our pixel-level experiments start by training the M2F-SwinL [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] on Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] and Vistas [Neuhold et al.(2017)Neuhold, Ollmann, Rota Bulo, and Kontschieder] with the Cityscapes taxonomy. We extend the mask-wide classifier to K+2 classes and fine-tune on mixed-content images. We compose mixed-content images by pasting negative content (ADE20k [Zhou et al.(2019)Zhou, Zhao, Puig, Xiao, Fidler, Barriuso, and Torralba] instances or flow samples) into inlier images. More implementation details are in the Appendix.
Segmentation of road scenes. Table 1 compares UNO with the related work on Fishyscapes and SMIYC. We observe that mask-level recognition methods outperform earlier works [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Vojir et al.(2021)Vojir, Šipka, Aljundi, Chumerin, Reino, and Matas]. When training with real negatives, our method outperforms all previous work on Fishyscapes and SMIYC AnomalyTrack while performing within variance of the best method on SMIYC ObstacleTrack. When training without real negatives, our method outperforms all previous work on SMIYC and Fishyscapes Static by a large margin while achieving second best AP on Fishyscapes L&F which appears to be a very challenging test benchmark for mask-level recognition methods because it includes small anomalies.
@@
NFlowJS [Grcić et al.(2021)Grcić, Bevandić, and Šegvić] 39.4 9.0 52.1 15.4 - - - - 77.4
SynDHybrid [Grcić and Šegvić(2023)] 51.8 11.5 54.7 15.5 - - - - 79.9
@@
Table 2 compares our method with other mask-level methods on validation subsets of Fishyscapes and RoadAnomaly [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann]. Our method outperforms all related approaches on all datasets while only having 0.5 pp worse $\text{FPR}_{95}$ than RbA on RoadAnomaly.
@@
Image classification. Table 3 shows the image-wide performance of UNO on the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. We compare UNO with training methods with and without the use of real negative data. When training with real negatives UNO consistently outperforms all baselines on CIFAR-10 and near OOD Imagenet-200 while attaining the best AUROC and the second-best FPR on the far OOD Imagenet-200. When training without real negative data UNO accomplishes competitive performance on both benchmarks. Note that UNO is indeed a training regularisation method, but with data as the regularisation source contrary to specially designed training algorithms. Synthetic UNO outperforms even methods trained with real data, specifically, MixOE, MCD and UDG on CIFAR-10 and UDG on ImageNet-200. It is interesting that UNO trained with synthetic negatives outperforms UNO trained with real negatives by a significant margin on the far OOD ImageNet-200. Finally, UNO preserves the inlier classification performance.
@@
Validating components of UNO score. Table 4 validates the contribution of the UNO components. In the case of real negatives, both components attain competitive results while UNO significantly outperforms both of them. In the case of synthetic negatives, uncertainty outperforms negative objectness by a wide margin. This does not come as a surprise since the negative objectness is inferred from synthetic training negatives. Still, UNO greatly benefits from the ensemble of the two scores.
@@
Figure 5: Qualitative experiments on Fishyscapes L&F val. Top row shows the input image and the three outlier scores. Bottom row shows anomaly detection maps after thresholding at TPR=95%. UNO significantly reduces false-positive responses.
@@ 6 Conclusion
Visual recognition models behave unpredictably in presence of outliers. A common strategy to detect outliers is to fine-tune the closed-set classifier in order to increase uncertainty in negative data. We take a different approach and disentangle the prediction uncertainty from the recognition of outliers. This allows us to cast outlier detection as an ensemble of in-distribution uncertainty and the posterior of the negative class, which we term negative objectness. The resulting UNO score manifests as a lightweight plug-in module that can extend an arbitrary classifier. The resulting model is then fine-tuned with real or synthetic negative data that imitate test outliers. UNO can be applied to per-pixel outlier detection by extending recent mask-level recognition approaches. There we first apply UNO to recover mask-wide outlier scores, and then propagate them to pixels according to mask assignments. UNO attains strong performance on recent image-wide and pixel-level outlier detection benchmarks with and without real negative data. Prominent future developments include improving the quality of synthetic negative data and extending our ensemble with generative predictions. Limitations of UNO are discussed in the Appendix.

### v4

@@ 4 Training UNO with and without real negative data
Every training minibatch contains an equal number of samples for all K+1 classes. Thus, we expose our model to significantly less negative data than the alternative approaches [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li, Yu and Aizawa(2019), Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu, Hendrycks and Gimpel(2017)], which utilize 2×2$\times$2 × more negatives than inliers. Even though the outlier exposure of our model is weak, it will still be biased towards detection of the test outliers that are similar to the training negatives. This issue can be circumvented by replacing the auxiliary negative dataset with synthetic negatives produced by a generative model [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Grcić et al.(2021b)Grcić, Bevandić, and Šegvić].
Synthetic negatives.
We generate model-specific synthetic negatives by a normalizing flow $f_{\psi}$ [Dinh et al.(2017)Dinh, Sohl-Dickstein, and Bengio, Kingma and Dhariwal(2018)] that we train alongside our discriminative classifier $f_{\theta}$. We train the flow to generate samples that resemble inliers, but at the same time give raise to uncertain predictions across the inlier classes according to the following loss [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic]:
@@
The first loss term $L_{\mathrm{mle}}(\psi)=\textbf{E}_{\mathbf{x^{+}}\in D_{\text{in}}}[-\ln p_{\psi}(\mathbf{x})]$ maximizes the likelihood of inlier samples, while the second loss term $L_{\text{jsd}}(\psi,\theta)=\text{JSD}(U,f_{\theta}(\mathbf{x}))$ minimizes the Jensen-Shannon divergence between the uniform distribution and the predictions over the K inlier classes. The two competing objectives settle down when the generative model produces samples at the border of the inlier manifold [Lee et al.(2018a)Lee, Lee, Lee, and Shin]. The above loss jointly optimizes $f_{\theta}$ and $f_{\psi}$ so that the discriminative model get aware of the negative class and thus preclude feature collapse [Lucas et al.(2019)Lucas, Shmelkov, Alahari, Schmid, and Verbeek]. Of course, the training also has to ensure that $f_{\theta}$ satisfies the primary discriminative loss.
However, a naive application of this recipe in the K+1-way context would not work as intended. In fact, the generative loss (6) and the discriminative loss (5) would simply be satisfied by confident classification of all generated samples into the negative class. This trivial solution would embed the generated samples into a compact region of the feature space, and make them useless as proxies for outlier detection. Thus, we propose a two-step optimization procedure. The first step jointly trains $f_{\psi}$, $f_{\theta}$, and the K inlier logits according to (6). The second step freezes the flow and fine-tunes the loss (5) in order to ensure recognition of generated negatives as the K+1-th class. Different from previous approaches [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic], the loss (6) serves only to promote generation of useful synthetic negatives, while we ensure the outlier recognition quality by subsequent fine-tuning according to the loss (5).
@@
Figure 4 shows our training approach in the dense prediction context. We paste negative training content atop the regular image to produce a mixed-content image. The mixed-content image is fed to the K+2-way classifier that optimizes the cross-entropy objective, similar as in the image-wide context. In the case of synthetic negatives, we jointly optimize the dense classifier and the normalizing flow.
Figure 4: Fine-tuning of dense classifier equipped with UNO. We paste negative training data (either real or synthetic) atop regular inlier images. The resulting mixed-content image is fed to the dense classifier that optimizes cross-entropy loss over K+2 classes.
@@ 5 Experiments
Experimental setup.
We evaluate the UNO performance on standard image-wide and pixel-level benchmarks. In the pixel-level setup, we consider Fishyscapes [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] and SMIYC [Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann], the two prominent benchmarks for road driving scenes. In the image-wide setup, we consider the CIFAR-10 and ImageNet-200 setups from the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. We report the standard evaluation metrics AUROC and FPR95 in image-wide, and AP and FPR95 in pixel-level experiments. Our pixel-level experiments start by training the M2F-SwinL [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] on Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] and Vistas [Neuhold et al.(2017)Neuhold, Ollmann, Rota Bulo, and Kontschieder] with the Cityscapes taxonomy. We extend the mask-wide classifier to K+2 classes and fine-tune on mixed-content images. We compose mixed-content images by pasting negative content (ADE20k [Zhou et al.(2019)Zhou, Zhao, Puig, Xiao, Fidler, Barriuso, and Torralba] instances or flow samples) into inlier images. More implementation details are in the Appendix. Our code is publicly available here [^1] https://github.com/matejgrcic/Open-set-M2F.
Segmentation of road scenes.
Table 1 compares UNO with the related work on Fishyscapes and SMIYC. We observe that mask-level recognition methods outperform earlier works [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Vojir et al.(2021)Vojir, Šipka, Aljundi, Chumerin, Reino, and Matas]. When training with real negatives, our method outperforms all previous work on Fishyscapes and SMIYC AnomalyTrack, while performing within variance of the best method on SMIYC ObstacleTrack. When training without real negatives, our method outperforms all previous work on SMIYC and Fishyscapes Static by a large margin, while achieving second best AP on Fishyscapes L&F. Interestingly, the inlier recognition performance is not affected by the UNO fine-tuning. The initial closed-set model attains 83.5% mIoU on Cityscapes. Appending UNO and fine-tuning the resulting open-set model with real negatives increases classification performance to 83.7%.
@@
NFlowJS [Grcić et al.(2021b)Grcić, Bevandić, and Šegvić] 39.4 9.0 52.1 15.4 - - - - 77.4
SynDHybrid [Grcić and Šegvić(2024a)] 51.8 11.5 54.7 15.5 - - - - 79.9
@@
Table 2 compares UNO with previous works that work with mask-level recognition on validation subsets of Fishyscapes and RoadAnomaly [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann]. UNO outperforms previous approaches across datasets and metrics with the single exception of 0.5 pp worse $\text{FPR}_{95}$ than RbA on RoadAnomaly.
@@
Image classification.
Table 3 shows the image-wide performance of UNO on the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. We compare UNO with training methods with and without the use of real negative data. When training with real negatives, UNO consistently outperforms all baselines on CIFAR-10 and near OOD Imagenet-200, while attaining the best AUROC and the second-best FPR on the far OOD Imagenet-200. When training without real negative data, UNO accomplishes competitive performance on both benchmarks. Synthetic UNO outperforms even methods trained with real data, specifically, MixOE, MCD and UDG on CIFAR-10, and UDG on ImageNet-200. It is interesting that UNO trained with synthetic negatives outperforms UNO trained with real negatives by a significant margin on the far OOD ImageNet-200. Finally, UNO preserves the inlier classification performance.
@@
Validating components of UNO score.
Table 4 validates the contribution of the UNO components. In the case of real negatives, both components attain competitive results while UNO significantly outperforms both of them. In the case of synthetic negatives, uncertainty outperforms the negative objectness by a wide margin. This does not come as a surprise since the negative objectness is inferred from synthetic training negatives. Still, UNO again greatly benefits from the ensemble of the two scores.
@@
Figure 5: Qualitative experiments on Fishyscapes L&F val. Top row shows the input image and the three outlier scores. Bottom row shows anomaly detection maps after thresholding at TPR=95%. UNO significantly reduces the incidence of false-positive responses.
@@ 6 Conclusion
Visual recognition models behave unpredictably in presence of outliers. A common strategy to detect outliers is to fine-tune the closed-set classifier in order to increase uncertainty in negative data. We take a different approach and disentangle the prediction uncertainty from the recognition of outliers. This allows us to cast outlier detection as an ensemble of in-distribution uncertainty and the posterior of the negative class, which we term negative objectness. The resulting UNO score manifests as a lightweight plug-in module that can extend an arbitrary classifier. The resulting model is then fine-tuned with real negative data or their synthetic surrogates. UNO can be applied to per-pixel outlier detection by extending recent mask-level recognition approaches. There we first apply UNO to recover mask-wide outlier scores, and then propagate them to pixels according to mask assignments. UNO attains strong performance on recent image-wide and pixel-level outlier detection benchmarks with and without real negative data. Prominent future developments include improving the quality of synthetic negative data and extending our ensemble with generative predictions.
7 Limitations
Auxiliary training data. UNO relies on negative training data to deliver strong OOD detection performance. While standard applications offer plenty of real negative data, this might not be the case in domain-specific applications such as remote sensing and medical imaging. Still, we show that UNO can deliver strong performance when trained on synthetic negatives.
Confidence calibration. Our K+1-way classifier could be inadequately calibrated which would introduce bias into our uncertainty estimates. An interesting direction for future work is formulation of a better uncertainty estimate.